OpenStudy (anonymous):

What is the area of sector CED when DE = 15 yd? 125π yd2 10π yd2 18.75π yd2 225π yd2

RhondaSommer (rhondasommer):

A=πr^2 is area for circle

RhondaSommer (rhondasommer):

so \[15^2\] is what

OpenStudy (anonymous):

225

RhondaSommer (rhondasommer):

so what is your answer?

OpenStudy (anonymous):

225π yd2

RhondaSommer (rhondasommer):

yup :)

OpenStudy (anonymous):

thank you

RhondaSommer (rhondasommer):

no problem

RhondaSommer (rhondasommer):

medal me :)

RhondaSommer (rhondasommer):

and close the problem

RhondaSommer (rhondasommer):

*please*

OpenStudy (anonymous):

okay

OpenStudy (wolf1728):

225 PI square yards is wrong area of a circle = PI*radius*radius area of a circle = PI*15*15 area of a circle = PI * 225 IF we needed the area for the entire circle. We need the SECTOR area which is 30 degrees of the entire circle. SO the SECTOR area = 30/360 degrees of the circle or (1/12) of the area SECTOR area = PI*225*(1/12) SECTOR area = PI*18.75 The answer is C

OpenStudy (anonymous):

okay thank you so much